OpenStudy (anonymous):

Solve for y , where y is a real number. √y+11= √7y-14 (If there is more than one solution, separate them with commas.)

OpenStudy (anonymous):

25/6

OpenStudy (anonymous):

if the radical is over the whole thing.

OpenStudy (anonymous):

yes satellite73

OpenStudy (jit4won):

\[\sqrt{y}=\sqrt{7}y-25 \]

OpenStudy (anonymous):

\[\sqrt{y+11}=\sqrt{7y-14}\]

OpenStudy (anonymous):

l radical per equation

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

\[y+11=7y-14\] \[25=6y\] \[y=\frac{25}{6}\]

OpenStudy (amistre64):

the strategic use of ()s greatly inhance our ability to decipher the questions :)

OpenStudy (anonymous):

hmmm then just ignore that part if it makes it easier for you lol

OpenStudy (amistre64):

√(y+11)= √(7y-14) ; square the sides y +11 = 7y -14 25 = 6y y = 25/6